OpenStudy (anonymous):

The sum of two consecutive integers is at least 52. What is the least possible pair of integers? A. 27 and 28 B. 25 and 26 C. 24 and 25 D. 26 and 27 @geekfromthefutur

OpenStudy (geekfromthefutur):

If we let the first number be A and the second number B, we would write A+ B > 52, we also know that B = A + 1 so we plug that back in A + A + 1 > 52, solve this and round up to the nearest integer to get A, then plug back in to B = A + 1 to get ___.

OpenStudy (anonymous):

B. ?

OpenStudy (geekfromthefutur):

Ding ding

OpenStudy (geekfromthefutur):

Your correct ^~^

OpenStudy (mathstudent55):

B is incorrect. 25 + 26 = 51. 51 is not at least 52. It is less than 52. @geekfromthefutur explained what to do. A + A + 1 > 52 2A + 1 > 52 2A >51 A > 25.5 \(A \ge 26\) The smaller of the two numbers has to be at least 26.

OpenStudy (geekfromthefutur):

So its D

OpenStudy (mathstudent55):

@geekfromthefutur In your own words: Ding ding